Put-Call Parity The current price of a stock is $35, and the annual risk-free rate is 3%. A call option with a strike price of $31 and with 1 year until expiration has a current value of $6.60. What is the value of a put option written on the stock with the same exercise price and expiration date as the call option? Do not round intermediate calculations. Round your answer to the nearest cent.

How do you calculate the negative exponent and how do you calculate using excel? Thank you.
Answer #1

According to put-call parity;

C + PV(x) = P + S

P = C + PV(x) - S

= $6.60 + [$31 x e-0.03] - $35 = $6.60 + $30.08 - $35 = $1.68